OpenStudy (anonymous):

Express the solution of y"+w^2*y=g(t); y(0)=0, y'(0)=1 in terms of a convolution integral.

OpenStudy (loser66):

@ybarrap

OpenStudy (ybarrap):

Ok, you'll need to take Laplace transform to solve. Here's an example: http://tutorial.math.lamar.edu/Classes/DE/ConvolutionIntegrals.aspx

OpenStudy (loser66):

you got (s^2+w^2)Y(s) =G(s)+1, right?

OpenStudy (anonymous):

What's after s^2*Y(s)-sy(0)-y'(0)+?

OpenStudy (loser66):

so, \[Y(s)=\frac{G(s)}{s^2+w^2}+\frac{1}{s^2+w^2}\]got this part? stubborn girl?

OpenStudy (loser66):

ok, let back to the beginning.

OpenStudy (anonymous):

But how did you get there? I only got s^2*Y(s)-sy(0)-y'(0)?

OpenStudy (loser66):